# Revisiting the Berkeley Admissions data: Statistical Tests for Causal Hypotheses

### v1

@@ 3.1 Fairness Notions
So, if the outcome of the test is that conditional independence cannot be rejected ($M\in H^{0}_{\text{no-cf-obs}}$), then due to the aforementioned observations, we cannot conclude that the underlying causal model belongs to the causal null hypothesis of the interventional or counterfactual or graphical fairness notions, i.e., our conclusion is that fairness is “undecidable”. However, if the outcome of the statistical test is that there is enough evidence in the data to reject conditional independence($M\notin H^{0}_{\text{no-cf-obs}}$), then we can conclude that the underlying causal model does not belong to the causal null hypothesis of any of the fairness notions, i.e., there is unfairness.
@@ 4 Berkeley Case: With Latent Confounding Between Department And Outcome
We now take a more careful causal modeling approach. Instead of starting from variables and reasoning about structural equations that we allow, we start with assuming that all structural equations exist. [^5] Since this allows for causal cycles, this would require using the framework of simple SCMs (Bongers et al., 2021). For the Berkeley example, Figure 3 shows a causal graph of an SCM that we start with. We now provide rationale for ruling out few structural equations. Based on chronology of events, we rule out those where $D$ directly affects $S$, where $A$ directly affects $D$ and where $A$ directly affects $S$. We rule out unobserved common causes of $S$ and $D$, and $S$ and $A$ since we model $S$ to be sex at birth. While latent selection bias might introduce bidirected edges (Chen et al., 2024) that are incident on $S$, we assume for now that there is no selection bias in the dataset. The resulting class of SCMs has structural equations of the form
$S$ $=f_{S}(U_{S}),$
$D$ $=f_{D}(S,U,U_{D}),$ (2)
$A$ $=f_{A}(S,D,U,U_{A}),$
where $U,U_{S},U_{D}$ and $U_{A}$ denote independent exogenous random variables. We define $\mathbb{M}_{\text{cf}}$ to be the family of models parameterized by the above structural equations and the exogenous distribution. Further, we define $\mathbb{M}_{\text{cf}+}=\left\{M\in\mathbb{M}_{\text{cf}}:\forall s,P_{M}(S=s)>0\right\}$. For $M\in\mathbb{M}_{\text{cf}}$ (and $\mathbb{M}_{\text{cf}+}$), the causal graph is a subgraph of the one shown in Figure 2(b).
Although we arrived at allowing confounding between department and outcome through a careful causal modeling approach, this is not a novel consideration. In particular, Kruskal (Fairley and Mosteller, 1977, Pg 128-129) demonstrated an example where the existence of a latent confounder, such as state of residence, can render Bickel et al. (1975)’s analysis incorrect. Other natural latent confounders include, for example, level of department-specific technical skills that influence both the department choice of an applicant and the admissions outcome.
@@
$H^{0}_{\text{cf-inter}+}$$\triangleq\left\{M\in\mathbb{M}_{\text{cf}+}:\forall d,s,P_{M}{\left({A=1\mid\text{do}(S=s),\text{do}(D=d)}\right)}=P_{M}{\left({A=1\mid\text{do}(D=d)}\right)}\right\},$
$H^{0}_{\text{cf-ctrf}+}$$\triangleq\left\{M\in\mathbb{M}_{\text{cf}+}:\forall d,s,P_{M}(A^{\text{do}(S=s,D=d)}=A^{\text{do}(D=d)})=1\right\},$
@@ 4.1 Graphical Notion and the Instrumental Variable (IV) Inequalities
Since the IV inequalities are only necessary conditions, an arbitrary distribution on $X,Y,Z$ that satisfies the IV inequality does not necessarily imply that it is an entailed distribution of a model from the IV model class. Bonet (2001) showed that for the binary instrument, treatment and effect case, the IV inequalities are also sufficient conditions. In Theorem 17, we show that for the case where the instrument and outcome are binary and the treatment is discrete-valued with finite support, any distribution that satisfies the IV inequality is also entailed by some causal model from the IV model class. To the best of our knowledge, Theorem 17 is a novel result. We defer the proof to Section D.1.
@@
For the Berkeley admissions case, assuming for now that the true observational distribution over $S,D,A$ is known, the observational distribution satisfying the IV inequalities implies that there exists a causal explanation (model) where the directed edge $S\rightarrow A$ is absent, i.e., given that $P(A,D,S)\in{\cal P}_{\text{IV}+}$, there exists $M\in\mathbb{M}_{\text{IV}+}$ such that $P_{M}(A,D,S)=P(A,D,S)$. On the other hand, the observational distribution violating the IV inequalities does not necessarily imply that the edge $S\rightarrow A$ is present since the IV model class, $\mathbb{M}_{\text{IV}+}$, is only a subset of all the models that do not contain the edge $S\rightarrow A$ in the causal graph. For example, the existence of latent confounding between $S$ and $A$ in a model $M$ may result in $M\notin\mathbb{M}_{\text{IV}+}$, even though $G{\left({M}\right)}$ does not necessarily contain the directed edge $S\rightarrow A$. However, the causal modeling assumption that defined $\mathbb{M}_{\text{cf}+}$ rules out latent confounding between $S$ and $A$. Therefore, given our modeling assumptions, $H^{0}_{\text{cf-graph}+}=\mathbb{M}_{\text{IV}+}$, and in turn, we conclude that violating the IV inequalities implies that $M\in\mathbb{M}_{\text{cf}+}\backslash H^{0}_{\text{cf-graph}+}$. As in the previous section, it is possible that causal models that lie outside $H^{0}_{\text{cf-graph}+}$ (“unfair” models) induce observational distributions that lie in ${\cal P}_{\text{IV}+}$, i.e., satisfy the IV inequalities. Therefore, satisfying the IV inequalities is not conclusive evidence that the data-generating mechanism is fair, i.e., our conclusion should be that fairness is undecidable. In Section 5 we introduce a Bayesian test for the IV inequalities.
@@ 4.2 Bounds on Interventional Notion of Fairness
For $M\in\mathbb{M}_{\text{cf}+}$, the interventional notion of fairness is the CDE, which is not identifiable in our case. By a response-function parameterization (Balke, 1995; Balke and Pearl, 1997) of $M\in\mathbb{M}_{\text{cf}+}$, we can express the interventional distributions in Definition 16 as a linear function of response variables. Further, the observational distribution is also expressed as a linear function of the response variables. Using the symbolic linear programming approach of Balke (1995), we obtain upper and lower bounds in terms of the observational distribution, specifically, $P_{M}{\left({A,D\mid S}\right)}$. Indeed, Cai et al. (2008) express the same bounds which we reproduce below. The CDE given by
$P_{M}{\left({A=1\mid\text{do}(S=1),\text{do}(D=d)}\right)}-P_{M}{\left({A=1\mid\text{do}(S=0),\text{do}(D=d)}\right)},$
@@
$\left[\Pr\left(A=1,D=d\mid S=1\right)+\Pr\left(A=0,D=d\mid S=0\right)-1,\right.$
$\left.1-\Pr\left(A=0,D=d\mid S=1\right)-\Pr\left(A=1,D=d\mid S=0\right)\right].$
For the interventional notion of fairness, the CDE must be 00 for all $d$. By setting the lower bound to be at most 00 and the upper bound to be at least 00, we recover the IV inequalities in (3). While Cai et al. (2008) do not point out the connection to the IV inequalities, they find it “remarkable that we [they] get such a simple formula, consisting of only one additive expression in the lower bound and one additive expression in the upper bound". In the next subsection, we show that the connection to the IV inequalities is not a coincidence.
@@ 4.3 A Family of Equivalent Tests
The graphical and interventional fairness notions end up imposing identical constraints on the observational distribution. However, note that $H^{0}_{\text{cf-inter}+}\supseteq H^{0}_{\text{cf-graph}+}$. In fact, we prove in Section D.2 that $H^{0}_{\text{cf-graph}+}=H^{0}_{\text{cf-ctrf}+}\subset H^{0}_{\text{cf-inter}+}$. Models in $H^{0}_{\text{cf-inter}+}\backslash H^{0}_{\text{cf-graph}+}$ (Example 14) are such that the edge $S\rightarrow A$ exists in the causal graph and yet, the interventional queries in Definition 16 are equal. Given that the null hypothesis of the interventional fairness notion is a strict superset of that of the graphical fairness notion, we might expect the same relation to hold in the resulting set of observational distributions for these hypotheses, thus giving us potentially different tests. In contrast, like in Section 3, we show that the corresponding sets of observational distributions resulting from models in $H^{0}_{\text{cf-inter}+}$, $H^{0}_{\text{cf-ctrf}+}$, $H^{0}_{\text{cf-graph}+}$ are identical. Section D.3 contains the proof.
@@ 4.4 Comparison With Existing Fairness Notions
The utility of considering statistical tests is that we can now compare different fairness notions for a particular case with respect to the same causal modeling assumptions. In this section, we consider the three existing counterfactual fairness notions, namely the NDE (Nabi and Shpitser, 2018; Chiappa, 2019), and the counterfactual and path-dependent counterfactual fairness notions in Kusner et al. (2017).
For NDE, Kaufman et al. (2005) obtain bounds for the all-binary setting. Using these bounds, we obtain a strictly weaker test than the IV inequalities. [^7] Intuitively, the reason is the same as in Section 3; the NDE averages over departments, and a positive bias in one department may cancel out against a similarly strong negative bias in another department. Hence, vanishing NDE does not imply that each department takes fair decisions. We show in Section E.2.2 that the counterfactual notion of fairness of Kusner et al. (2017) implies demographic parity even when confounding is allowed. In Section E.2.3 we show that testing the path-dependent counterfactual fairness notion of Kusner et al. (2017) is equivalent to testing the IV inequalities.
@@ 5 Bayesian Testing Procedure
We start with proposing a Bayesian test for IV inequalities using finite data. Although Ramsahai (2008) proposed a frequentist test and derived the distribution of the likelihood ratio, providing a means to obtain a p-value for the case of a binary treatment, it is unclear what the test would be for our case where the “treatment” is not binary. In addition, Wang et al. (2017) provide a frequentist test that involves multiple one-sided independence tests. In contrast, the Bayesian test has straightforward extensions to other hypotheses that are expressed in terms of the observational distribution, including bounds on unidentifiable causal queries.
@@
Results on Berkeley admission data: We use the UCBAdmissions (R Core Team, 2023) dataset from R that contains counts for each sex-department-admissions outcome tuple for the $6$ largest departments. Therefore, $|{\cal X}|=6$, $|{\cal Y}|=2,|{\cal Z}|=2$. For parameters
$\theta={\left({P(d,a,s):s\in{\cal X}_{S},d\in{\cal X}_{D},a\in{\cal X}_{A}}\right)},$
we choose a flat Dirichlet prior over $\Delta$ giving us $\pi(\theta|\mathbb{M}_{i})=c_{i}\text{Dir}{\left({1,1,\cdots,1}\right)}\bm{1}\left[\theta\in\mathbb{M}_{i}\right]$ where $c_{i}$ is a normalizing constant. The counts from the data are used to obtain the posterior, $P(\theta\mid R_{1},R_{2},\cdots,R_{m})$ which is also a truncated Dirichlet distribution. Using $n=10^{6}$ samples, we observe no violations of the IV inequality. Therefore, the confidence interval for the posterior probability of the Berkeley data satisfying the IV inequalities is $\left[1-3.69\times 10^{-6},1\right]$. As mentioned in Section 4.1 satisfying the IV inequalities implies that fairness is undecidable. In Section F, we carry out a sensitivity analysis by varying the chosen prior. We also report results on a different dataset from Bol (2023) that investigates sex-based discrimination in awarding cum-laude distinctions to graduate students.
In addition to the Bayesian test, the maximum likelihood (ML) estimator satisfies the IV inequalities, implying that there isn’t enough evidence to reject the null hypothesis when doing a likelihood ratio test. An implementation of Wang et al. (2017) for the Berkeley dataset also does not reject the null hypothesis (see Section F for details).
@@ 6 Discussion
The Berkeley admissions case is a canonical example in the causal fairness literature. Bickel et al. (1975) reached the conclusion of rejecting unfairness while making unrealistic assumptions of no unobserved confounding. We take the next step of analyzing the Berkeley dataset while allowing for unobserved confounding which results in a different conclusion; since there is very strong evidence that the data satisfies the IV inequalities, the conclusion regarding fairness, from the available data, is that it is undecidable.
@@
While our analysis was centered around the Berkeley case, there are multiple aspects that generalize—a) The family of causal models we consider can be thought of as a mediator with a confounder between mediator and outcome, which is common in mediation analysis. b) The approach of fairness notions being causal hypotheses, with respect to the class of models defined by modeling assumptions, that need to be translated into statistical tests to be useful in practice. c) The observation that for the case of inequality constraints on observational data, a straightforward Bayesian testing procedure is available.
@@
Selection Bias: UCBAdmissions dataset only has data from the $6$ largest departments as opposed to $85$ in Bickel et al. (1975). Also, the fraction of female students is significantly smaller than the fraction of male students. There could possibly be other sources of latent selection that alter the causal model resulting in violating the assumptions of, for instance, absence of bidirected edge in the causal graph between $S$ and $A$. Since allowing for selection bias enlarges the model class $\mathbb{M}_{\text{cf}+}$, given that the data satisfies IV inequalities, we conclude that allowing for selection bias will not change our conclusion. We leave a deeper analysis that takes selection bias into account as future work.
7 Acknowledgements
This work was supported by Booking.com.

### v2

@@ 3.1 Fairness Notions
So, if the outcome of the test is that conditional independence cannot be rejected ($M\in H^{0}_{\text{no-cf-obs}}$), then due to the aforementioned observations, we cannot conclude that the underlying causal model belongs to the causal null hypothesis of the interventional or counterfactual or graphical fairness notions, i.e., our conclusion is that fairness is “undecidable” with respect to these notions. However, if the outcome of the statistical test is that there is enough evidence in the data to reject conditional independence ($M\notin H^{0}_{\text{no-cf-obs}}$), then we can conclude that the underlying causal model does not belong to the causal null hypothesis of any of the fairness notions. If we rule out the existence of any latent, unprotected mediator between $S$ and $A$, we can conclude that the data generating mechanism is unfair.
@@ 4 Berkeley Case: With Latent Confounding Between Department And Outcome
We now take a more careful causal modeling approach. Instead of starting from variables and reasoning about structural equations that we allow, we start with assuming that all structural equations exist. [^5] Since this allows for causal cycles, this would require using e.g. the framework of simple SCMs [Bongers et al., 2021]. For the Berkeley example, Figure 3 shows a causal graph of an SCM that we start with. We now provide rationale for ruling out few structural equations. Based on chronology of events, we rule out those where $D$ directly affects $S$, where $A$ directly affects $D$ and where $A$ directly affects $S$. We rule out unobserved common causes of $S$ and $D$, and $S$ and $A$ since we model $S$ to be sex at birth. While latent selection bias might introduce bidirected edges [Chen et al., 2024] that are incident on $S$, we assume for now that there is no selection bias in the dataset. The resulting class of SCMs has structural equations of the form $S=f_{S}(U_{S}),D=f_{D}(S,U,U_{D}),A=f_{A}(S,D,U,U_{A})$ where $U,U_{S},U_{D}$ and $U_{A}$ denote independent exogenous random variables. We define $\mathbb{M}_{\text{cf}}$ to be the family of models parameterized by the above structural equations and the exogenous distribution. Further, we define $\mathbb{M}_{\text{cf}+}=\left\{M\in\mathbb{M}_{\text{cf}}:\forall s,P_{M}(S=s)>0\right\}$. For $M\in\mathbb{M}_{\text{cf}}$ (and $\mathbb{M}_{\text{cf}+}$), the causal graph is a subgraph of the one shown in Figure 2(b).
Although we arrived at allowing confounding between department and outcome through a careful causal modeling approach, this is not a novel consideration. In particular, Kruskal [Fairley and Mosteller, 1977, Pg 128-129] demonstrated an example where the existence of a latent confounder, such as state of residence, can render Bickel et al. [1975]’s analysis incorrect. Other natural latent confounders include, for example, level of department-specific technical skills that influence both the department choice of an applicant and the admissions outcome.
@@
$H^{0}_{\text{cf-inter}+}$$\triangleq\left\{M\in\mathbb{M}_{\text{cf}+}:\forall d,s,\right.$
$\left.P_{M}{\left({A=1\mid\text{do}(S=s),\text{do}(D=d)}\right)}\right.$
$\left.=P_{M}{\left({A=1\mid\text{do}(D=d)}\right)}\right\},$
$H^{0}_{\text{cf-ctrf}+}$$\triangleq\left\{M\in\mathbb{M}_{\text{cf}+}:\forall d,s,\right.$
$\left.P_{M}(A^{\text{do}(S=s,D=d)}=A^{\text{do}(D=d)})=1\right\},$
@@ 4.1 Graphical Notion and the Instrumental Variable (IV) Inequalities
Since the IV inequalities are only necessary conditions, an arbitrary distribution on $X,Y,Z$ that satisfies the IV inequality does not necessarily imply that it is an entailed distribution of a model from the IV model class. Bonet [2001] showed that for the binary instrument, treatment and effect case, the IV inequalities are also sufficient conditions. In Theorem 17, we show that for the case where the instrument and outcome are binary and the treatment is discrete-valued with finite support, any distribution that satisfies the IV inequality is also entailed by some causal model from the IV model class. To the best of our knowledge, Theorem 17 is a novel result. We defer the proof to Section D.1.
@@
Let $X,Y,Z$ be discrete random variables defined on ${\cal X},{\cal Y},{\cal Z}$ respectively, with $|{\cal X}|=n\geq 2,|{\cal Y}|=2,|{\cal Z}|=2$. Define ${\cal P}_{\mathbb{M}_{\text{IV}+}}\triangleq\left\{P_{M}(X,Y,Z):M\in\mathbb{M}_{\text{IV}+}\right\}$ and the set of joint distributions that satisfy the IV inequalities as
@@
[…]
Then ${\cal P}_{\text{IV}+}={\cal P}_{\mathbb{M}_{\text{IV}+}}.$
For the Berkeley admissions case, the observational distribution satisfying the IV inequalities implies that there exists a causal explanation (model) where the directed edge $S\rightarrow A$ is absent, i.e., given that $P(A,D,S)\in{\cal P}_{\text{IV}+}$, there exists $M\in\mathbb{M}_{\text{IV}+}$ such that $P_{M}(A,D,S)=P(A,D,S)$. On the other hand, the observational distribution violating the IV inequalities does not necessarily imply that the edge $S\rightarrow A$ is present since the IV model class, $\mathbb{M}_{\text{IV}+}$, is only a subset of all the models that do not contain the edge $S\rightarrow A$ in the causal graph. For example, the existence of latent confounding between $S$ and $A$ in a model $M$ may result in $M\notin\mathbb{M}_{\text{IV}+}$, even though $G{\left({M}\right)}$ does not necessarily contain the directed edge $S\rightarrow A$. However, the causal modeling assumption that defined $\mathbb{M}_{\text{cf}+}$ rules out latent confounding between $S$ and $A$. Therefore, given our modeling assumptions, $H^{0}_{\text{cf-graph}+}=\mathbb{M}_{\text{IV}+}$, and in turn, we conclude that violating the IV inequalities implies that $M\in\mathbb{M}_{\text{cf}+}\backslash H^{0}_{\text{cf-graph}+}$. As in the previous section, it is possible that causal models that lie outside $H^{0}_{\text{cf-graph}+}$ (“unfair” models) induce observational distributions that lie in ${\cal P}_{\text{IV}+}$, i.e., satisfy the IV inequalities. Therefore, satisfying the IV inequalities is not conclusive evidence that the data-generating mechanism is fair, i.e., our conclusion should be that fairness is undecidable. In Section 5 we introduce a Bayesian test for the IV inequalities.
@@ 4.2 Bounds on Interventional Notion of Fairness
For $M\in\mathbb{M}_{\text{cf}+}$, the interventional notion of fairness is equivalent to a vanishing CDE, where the latter is not identifiable in our case. By a response-function parameterization [Balke, 1995, Balke and Pearl, 1997] of $M\in\mathbb{M}_{\text{cf}+}$, we can express the interventional distributions in Definition 16 as a linear function of response variables. Further, the observational distribution is also expressed as a linear function of the response variables. Using the symbolic linear programming approach of Balke [1995], we obtain upper and lower bounds in terms of the observational distribution, specifically, $P_{M}{\left({A,D\mid S}\right)}$. Indeed, Cai et al. [2008] express the same bounds which we reproduce below. The CDE
$P_{M}{\left({A=1\mid\text{do}(S=1),\text{do}(D=d)}\right)}$
$-P_{M}{\left({A=1\mid\text{do}(S=0),\text{do}(D=d)}\right)},$
@@
$\left[\Pr\left(A=1,d\mid S=1\right)+\Pr\left(A=0,d\mid S=0\right)-1,\right.$
$\left.1-\Pr\left(A=0,d\mid S=1\right)-\Pr\left(A=1,d\mid S=0\right)\right].$
For the interventional notion of fairness, the CDE must be 00 for all $d$. By setting the lower bound to be at most 00 and the upper bound to be at least 00, we recover the IV inequalities in (2). While Cai et al. [2008] do not point out the connection to the IV inequalities, they find it “remarkable that we [they] get such a simple formula, consisting of only one additive expression in the lower bound and one additive expression in the upper bound”. In the next subsection, we show that the connection to the IV inequalities is not a coincidence.
@@ 4.3 A Family of Equivalent Tests
The graphical and interventional fairness notions end up imposing identical constraints on the observational distribution. However, note that $H^{0}_{\text{cf-inter}+}\supseteq H^{0}_{\text{cf-graph}+}$. In fact, we prove in Section D.2 that $H^{0}_{\text{cf-graph}+}=H^{0}_{\text{cf-ctrf}+}\subset H^{0}_{\text{cf-inter}+}$. Models in $H^{0}_{\text{cf-inter}+}\backslash H^{0}_{\text{cf-graph}+}$ (Example 14 with $\varepsilon=\frac{1}{2}$) are such that the edge $S\rightarrow A$ exists in the causal graph and yet, the interventional queries in Definition 16 are equal. Given that the null hypothesis of the interventional fairness notion is a strict superset of that of the graphical fairness notion, we might expect the same relation to hold in the resulting set of observational distributions for these hypotheses, thus giving us potentially different tests. In contrast, like in Section 3, we show that the corresponding sets of observational distributions resulting from models in $H^{0}_{\text{cf-inter}+}$, $H^{0}_{\text{cf-ctrf}+}$, $H^{0}_{\text{cf-graph}+}$ are identical. Section D.3 contains the proof.
@@ 4.4 Comparison With Existing Fairness Notions
The utility of considering statistical tests is that we can now compare different fairness notions for a particular case with respect to the same causal modeling assumptions. In this section, we consider the three existing counterfactual fairness notions, namely the NDE [Nabi and Shpitser, 2018, Chiappa, 2019], and the counterfactual and path-dependent counterfactual fairness notions in Kusner et al. [2017].
For NDE, Kaufman et al. [2005] obtain bounds for the all-binary setting. Using these bounds, we obtain a strictly weaker test than the IV inequalities. [^7] Intuitively, the reason is the same as in Section 3; the NDE averages over departments, and a positive bias in one department may cancel out against a similarly strong negative bias in another department. Hence, vanishing NDE does not imply that each department takes fair decisions. We show in Section E.2.2 that the counterfactual notion of fairness of Kusner et al. [2017] implies demographic parity even when confounding is allowed. In Section E.2.3 we show that testing the path-dependent counterfactual fairness notion of Kusner et al. [2017] is equivalent to testing the IV inequalities.
@@ 5 Bayesian Testing Procedure
We start with proposing a Bayesian test for IV inequalities using finite data. Although Ramsahai [2008] proposed a frequentist test and derived the distribution of the likelihood ratio, providing a means to obtain a p-value for the case of a binary treatment, it is unclear what the test would be for our case where the “treatment” is not binary. In addition, Wang et al. [2017] provide a frequentist test that involves multiple one-sided independence tests. In contrast, the Bayesian test has straightforward extensions to other hypotheses that are expressed in terms of the observational distribution, including bounds on unidentifiable causal queries.
@@
Results on Berkeley admission data: We use the UCBAdmissions [R Core Team, 2023] dataset from R that contains counts for each sex-department-admissions outcome tuple for the $6$ largest departments. Therefore, $|{\cal X}|=6$, $|{\cal Y}|=2,|{\cal Z}|=2$. Since the data satisfies the positivity for sex, we can use the Bayesian model selection procedure. For parameters $\theta={\left({P(d,a,s):s\in{\cal X}_{S},d\in{\cal X}_{D},a\in{\cal X}_{A}}\right)}$,we choose a flat Dirichlet prior over $\Delta$ giving us $\pi(\theta|\mathbb{M}_{i})=c_{i}\text{Dir}{\left({1,1,\cdots,1}\right)}\bm{1}\left[\theta\in\mathbb{M}_{i}\right]$ where $c_{i}$ is a normalizing constant. The counts from the data are used to obtain the posterior, $P(\theta\mid R_{1},R_{2},\cdots,R_{m})$ which is also a truncated Dirichlet distribution. Using $n=10^{6}$ samples, we observe no violations of the IV inequalities. Therefore, the confidence interval for the posterior probability of the Berkeley data satisfying the IV inequalities is $\left[1-3.69\times 10^{-6},1\right]$. As mentioned in Section 4.1 satisfying the IV inequalities implies that fairness is undecidable. In Section F, we carry out a sensitivity analysis by varying the chosen prior. We also report results on a different dataset from Bol [2023] that investigates sex-based discrimination in awarding cum-laude distinctions to graduate students.
In addition to the Bayesian test, the maximum likelihood (ML) estimator satisfies the IV inequalities, implying that there isn’t enough evidence to reject the null hypothesis when doing a likelihood ratio test. An implementation of Wang et al. [2017] for the Berkeley dataset also does not reject the null hypothesis (see Section F for details). The code can be found at https://github.com/SourbhBh/BerkeleyCode.
@@ 6 Discussion
The Berkeley admissions case is a canonical example in the causal fairness literature. Bickel et al. [1975] reached the conclusion of there being no evidence to reject fairness, although under the unrealistic assumption of no unobserved confounding. When allowing for unobserved confounding, we arrived at a different conclusion: since there is very strong evidence that the data satisfies the IV inequalities, it is undecidable from the available data whether the admission procedure was fair or discriminated against sex.
@@
While our analysis was centered around the Berkeley case, there are multiple aspects that generalize—a) $\mathbb{M}_{\text{cf}}$ can be thought of as a mediator with a confounder between mediator and outcome, which is common in mediation analysis. b) The approach of fairness notions being causal hypotheses, with respect to the class of models defined by modeling assumptions, that need to be translated into statistical tests to be useful in practice. c) The observation that for the case of inequality constraints on observational data, a straightforward Bayesian testing procedure is available.
Generalization: Non-binary variables: While the Berkeley dataset had a binary protected attribute and a binary decision outcome, our approach can be generalized to non-binary variables. The response-function parametrization of $\mathbb{M}_{\text{IV}+}$ can be used to characterize the set of induced observed distributions of $X,Y$ and $Z$ as a convex polyhedral set. Using computer algebra, this polyhedral set can be characterized by linear inequality constraints even when the alphabet sizes are arbitrary. However, the number of linear inequality constraints quickly explodes for non-binary $Z$ (e.g. for binary $X$ and $Y$, and for $|{\cal Z}|=2,3,4$ and $5$, we get $12,48,160$ and $420$ inequalities). Nevertheless, these sets of inequality constraints can be tested using a Bayesian testing procedure akin to the one we outline in Section 5, thus giving a statistical test for all our fairness notions since the statement of Theorem 18 holds for any alphabet size.
Generalization: Relaxing unobserved confounding assumptions In Section 4 we made the assumption of no unobserved confounding between $S$ and any other variable. In Section D.4 we show that our main theorems, Theorem 17 and Theorem 18 hold even in the case of allowing for confounding between $S$ and $D$. If unobserved confounding is allowed between $S$ and $A$, the set of observational distributions induced by causal models where the direct effect between sex and admissions outcome is absent is the entire simplex which is the same as the set of observational distributions induced by causal models where the direct effect between sex and admissions outcome is present. Therefore, from observational data, it is not possible to distinguish between the presence/absence of a direct effect from sex to admissions outcome if we allow for confounding between sex and admissions outcome.
Unmeasured Mediators: Our conclusions and fairness notions are with respect to the measured variables. The presence of unmeasured mediators could change the interpretation of the results. For example, an unmeasured mediator that is not a ‘protected’ variable, such as choice of undergraduate department, would result in a direct effect of sex on admissions outcome in the marginalized causal model, but might still be considered ‘fair’. Therefore, even if the data does not satisfy the IV inequalities, we can only claim the existence of the direct effect from sex to admissions outcome and not whether there is unfairness. However, if the existence of latent unprotected mediators is ruled out, then we can consider the existence of a direct effect as ‘unfair’.
Undecidability: Our analysis can be viewed as a Bayesian model comparison between causal models (i) without a direct effect of $S$ on $A$, vs. (ii) with a direct effect of $S$ on $A$. If we only have access to observational data then causal models in (ii) result in a saturated model in the observed distribution space, i.e., the set of induced observational distributions is the entire simplex, whereas for the causal models in (i), the IV inequalities hold. Therefore, satisfying the IV inequalities implies that fairness is ‘undecidable’ since causal models in both (i) and (ii) satisfy the IV inequalities. Violating the IV inequalities, however, would imply that there is a direct effect of sex on admissions outcome. Only with the additional modeling assumption of there being no unprotected mediators between $S$ and $A$, could it then be concluded that the admissions process was ‘unfair’.
@@
Selection Bias: UCBAdmissions dataset only has data from the $6$ largest departments as opposed to $85$ in Bickel et al. [1975]. Also, the fraction of female students is significantly smaller than the fraction of male students. Hence, it is plausible that (latent) selection mechanisms alter the causal model resulting in violating the assumptions of, for instance, absence of a bidirected edge in the causal graph between $S$ and $A$ [Chen et al., 2024]. Since allowing for selection bias enlarges the model class $\mathbb{M}_{\text{cf}+}$, and given that the data satisfies the IV inequalities, we conclude that allowing for selection bias will not change our conclusion.
Acknowledgements.
This work was supported by Booking.com. We thank an anonymous reviewer for their feedback which helped us modulate some of our conclusions.